OpenStudy (anonymous):

How would I solve this problem exactly? 4.6 x 7.86=

OpenStudy (anonymous):

7.86 x4.60 ____ 6x6=36 carry the three _____ 6 6x8=48+three=51 carry the five ______ 16 6x7=42+5=47 ______ 4716 4x6=24 carry the two ______ 4716 4 4x8=32+two=34 carry the three _______ 4716 44 4x7=28+three+31 _______ 4716 3144 The spaces don't display correctly, however the 3144 is shifted one digit to the left of lining up with 4716, add together. _______ 36156 You had 3 decimal places in the two factors, 7.86 and 4.6, so go 3 places from the right and add the decimal. ________ 36.156

OpenStudy (anonymous):

I appreciate the break down like you did but maybe you could show me your work only please because I keep coming up with a different answer. thank you